The amount of rice cake mix to be used varies directly with the number of people who partakes of this food. Suppose that it takes 7 cups of mix to serve 10 people.
a.) Write the function expressing the number of cups of cake mix in relation to the number of people.
b.) How many cups of cake mix must be prepared for 50 people?
c.) About how many people can be served with 12 cups of cake mix?
These are my unsure answers:
For the letter b, i've tried solving it like this:
= 7/10=x/50
= 10x/10=350/10
= x=35
For letter c, I tried this:
= 12/7=x/10
= 7x/7=120/7
= x=17.14
So basically i don't understand letter a.

Answer by rothauserc(4718)

let y be the amount of cake mix
let x be the number of people
k is the constant of proportionality
then we have the following equation
y = kx
a) 7 = k10
k = 7/10 = .7
y = .7x
b) y = .7 * 50 = 35 cups of cake mix
c) 12 = .7x
x = 12/.7 = 17.14 people